OpenStudy (jenniferjuice):

Multiply -3x2(6x2 + 2x - 3) -18x2 - 6x + 9 -18x4 - 6x3 + 9x2 3x2 - x - 6 3x4 - x3 - 6x2

hartnn (hartnn):

distribute -3x^2 to each of the terms

hartnn (hartnn):

\(\Large -3x^2 \times 6x^2 = ... ?\)

hartnn (hartnn):

\(\Large -3x^2 \times 6x^2 =(-3\times6) x^{2+2}= ... ?\)

OpenStudy (jenniferjuice):

-18x4 - 6x3 + 9x2?

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

right

OpenStudy (jenniferjuice):

thank you :D

OpenStudy (jenniferjuice):

@hartnn

hartnn (hartnn):

welcome ^_^